OpenStudy (acacia):

Which is the maximum value for the vertices for the following: y≤2x−3y≤−13x+4 x≤9 f( x,y )=3x−2y Help?

OpenStudy (anonymous):

is that: y≤2x −3y≤−13x+4 x≤9 separately?

OpenStudy (anonymous):

is x and y greater than zero also? like a standard condition or are u just given x≤9?

OpenStudy (acacia):

I was just givin the x≤9

OpenStudy (acacia):

y≤2x-3 y≤−1/3x+4 sorry wrote it wrong

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i would first graph all three lines together on the same graph but graph them as y=2x-3 y=−x/3+4 x=9

OpenStudy (anonymous):

follow?

OpenStudy (acacia):

following

OpenStudy (anonymous):